# Reanimating Images using Neural Representations of Dynamic Stimuli

### v2

@@ Abstract
While computer vision models have made incredible strides in static image recognition, they still do not match human performance in tasks that require the understanding of complex, dynamic motion. This is notably true for real-world scenarios where embodied agents face complex and motion-rich environments. Our approach leverages state-of-the-art video diffusion models to decouple static image representation from motion generation, enabling us to utilize fMRI brain activity for a deeper understanding of human responses to dynamic visual stimuli. Conversely, we also demonstrate that information about the brain’s representation of motion can enhance the prediction of optical flow in artificial systems. Our novel approach leads to four main findings: (1) Visual motion, represented as fine-grained, object-level resolution optical flow, can be decoded from brain activity generated by participants viewing video stimuli; (2) Video encoders outperform image-based models in predicting video-driven brain activity; (3) Brain-decoded motion signals enable realistic video reanimation based only on the initial frame of the video; and (4) We extend prior work to achieve full video decoding from video-driven brain activity. This framework advances our understanding of how the brain represents spatial and temporal information in dynamic visual scenes. Our findings demonstrate the potential of combining brain imaging with video diffusion models for developing more robust and biologically-inspired computer vision systems. We show additional decoding and encoding examples on this site: https://sites.google.com/view/neural-dynamics/home.
@@ 1 Introduction
Figure 1: Encoding and decoding video motion using brain activity. (a) We demonstrate that fMRI brain activity can be reasonably predicted (as measured with a Pearson coefficient) from off-the-shelf video encoders (VideoMAE [1]) extracted from the viewed video. We also examine the converse direction; generating video by decoding brain activity. (b) Many existing video diffusion models (e.g., SVD [2]) generate a video by animating an initial frame. (c) This suggests one avenue for “brain-to-video” generation; fine-tune such diffusers to condition on fMRI input (e.g., MindVideo [3]). (d) Instead, we explicitly decouple the task of image and motion generation from brain activity. Given an initial video frame (which could be decoded from brain activity as in (c)) and fMRI input, we train a network to predict optical flow. We then animate the initial frame by feeding the predicted flow into an off-the-shelf motion-conditioned diffusion model (e.g., DragNUWA[4]). Our disentangled pipeline produces far more accurate brain-conditioned motion decodings than the prior art of (b) or (c). (e) We show the ground truth optical flow and video.
@@
Human vision excels at understanding complex, dynamic scenes, while computer vision offers powerful tools for data processing and analysis. By combining these complementary strengths, we can develop more robust and adaptable vision systems [5]. This work explores how insights from human brain activity, specifically in response to dynamic visual stimuli, can enhance the ability of computer vision systems to understand complex motion. Critically, as inherently embodied systems, our brains have evolved to process and represent the daily experience of naturally motion-rich environments in ways that are both semantically grounded and behaviorally-relevant. For example, a video of a person walking affords information not only about the visual features of the walker, but also about motion patterns, context, intent, and relationships between the different elements in the scene (e.g., sudden increase in gait implies urgency). This interplay between spatial and temporal information is impossible to capture from static images; accordingly, humans excel at inferring such high-level information from scene dynamics [6].
@@
While recent work has made significant strides in understanding the neural representation of visual information using computer vision models [7, 8] there remains a notable gap in our understanding of how the brain processes dynamic visual stimuli. Although some studies have begun to investigate the neural representation of dynamic visual stimuli [3, 9, 10], there is still a vast opportunity to leverage modern computer vision techniques, particularly video-based models. Second there has been little work demonstrating how insights from human vision can be leveraged to build better computer vision systems. Here we address both challenges, demonstrating that state-of-the-art video foundation models can be used to robustly predict human brain responses to dynamic visual inputs and that brain-decoded motion signals can help to reanimate videos based only on the first frame of the video. Critically, our framework outperforms prior work for both problems, achieving lower end point error (EPE) for motion decoding based on brain responses as compared to [3] and better motion prediction based on initial static frames with brain data than without brain data.
@@
In this paper, we present a framework for analyzing dynamic visual stimuli that explicitly separates the modeling of static and motion-related image representations as they are processed in the human brain. By decoupling these elements, our framework offers a targeted approach to understanding the distinct neural pathways used to process and represent dynamic visual information in real-world scenes.
While an important question in its own right, better specified models of dynamic visual processing are also critical for developing effective models of ecologically critical domains such as social interaction, action representation, and dynamic causal structure – none of which are easily extracted from static images. For example, social roles are are often inferred through the relative physical motions of different actors [6]. Similarly, many different actions appear the same in static images, for instance, opening versus closing a door, or standing up versus sitting down. Finally, learning about functional world knowledge (e.g., tool use or context-appropriate physical interactions) is primarily acquired through observation and imitation of dynamic events [11]. For these and many other domains, adding motion removes ambiguity.
@@
Immediate Contributions. In this context, our project makes several immediate contributions. First, we demonstrate that visual motion information in the form of optical flow can be predicted (or decoded) from brain activity as measured by fMRI. Second, we show that this predicted motion can be used to realistically animate static images using a motion-conditioned video diffusion model in which the motion is driven by brain activity. Moreover, we demonstrate that brain activity can produce more reliable animations than image-conditioned video diffusion models trained on massively-large datasets. Third, we apply our method to existing methods that decode the viewed video by using their initial frame to do full video decoding from video-driven brain activity and find that we more accurately predict the viewed motion than existing methods. Finally, we show prediction in the reverse direction: recent video encoders (such as VideoMAE and VC-1 [1, 12]) can more effectively predict fMRI brain activity from video imagery as compared to image encoders. Overall these results demonstrate the effectiveness of our approach as a novel, extensible framework for interpreting how the human brain processes dynamic visual information.
@@
Larger Objectives. Our project also has several larger objectives. First, developing high-performing models for motion decoding based on neural data (Figure 1d) is an important precursor to building effective encoding models [13]. Second, the finding that VideoMAE is able to accurately predict brain responses (Figure 1a) is an important step towards building more general and informative video encoding models for brain prediction [8]). Encoding models are also important components in comparisons between model-based representations and empirical neural data, which can reveal key insights into which visual features the brain prioritizes and how they align with computational processing in interpretable artificial neural networks [14]. Third, we apply our method to existing methods which decode the viewed video to do full video decoding from video-driven brain activity [3] and find that we more accurately predict the viewed motion than existing methods. Fourth, although comparison points are limited due to the underexploration of this research area, our finding that neural data improves video prediction performance better than other extant models is a simple demonstration of the point that latent information encoded in brain data may be useful for improving the performance of computer vision systems [15]. This approach may prove to be especially powerful for domains that are inherently anchored in human behavior, for example, dynamic social interactions.
@@ 2 Related Work
Static Visual Stimuli Decoding from fMRI. Reconstructing visual images from human brain activity, such as that measured by functional magnetic resonance imaging (fMRI), has significantly advanced through the use of deep learning models trained on extensive datasets of natural images [7, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31]. Recent methods leverage deep generative models to analyze and decode brain representations [32, 33, 3, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58]. This progress is largely due to the availability of large, publicly accessible fMRI datasets with well-annotated stimuli, such as BOLD5000 [59] and the Natural Scenes Dataset (NSD) [60], along with multimodal deep generative models that have been pre-trained on very large datasets, such as Stable Diffusion [61].
@@
Our fMRI-based approach offers a whole-brain perspective compared to single-neuron recordings, which are often limited in spatial coverage [65, 66, 67, 62, 63, 68]. Despite the challenges of fMRI’s relatively slow temporal resolution, our results demonstrate successful decoding of motion information from complex, naturalistic stimuli at the object level. This suggests that fMRI, despite its limitations, can still provide valuable insights into the neural representation of dynamic information.
Figure 2: Pipeline for motion decoding and video generation. (a) The model takes in neural data and image features, from DINOv2 [69], extracted from the initial frame to predict consecutive future dense optical flow fields. We then mask out salient objects, using FlowSAM [70], to obtain the masked object flow. Snowflakes denote frozen models, while flames denote actively trained models. (b) We realistically animate the initial frame using our predicted motion and a motion-conditioned video diffusion model, DragNUWA [4].
Concurrent to motion decoding research, there is a parallel branch of research focusing on decoding the raw video from fMRI activity [71, 72, 3, 73, 74, 75]. Methods like MindVideo [3] train to align temporal semantics from fMRI and generate frames across time. Notably, our approach explicitly disentangles the representations of motion information (represented as optical flow) from static images in fMRI signals. This allows us to differentiate our present work from previous work by enabling interpretability. Specifically, we isolate and focus on decoding motion (as optical flow) from fMRI signals, as opposed to more difficult to interpret mixture of static image features and motion. This disentanglement allows us to learn an explicit representation of motion which can be directly compared to groundtruth and be more directly related to spatially localized neural mechanisms, as opposed to an entangled mixture of static image and dynamic features.
@@ 3.1 Motion Prediction
Our primary objective is to extract the information relevant to visual motion in order to predict the optical flow in videos observed by participants. To enhance interpretability, we’ve architecturally separated static image features from dynamic features. By conditioning the model on the initial frame alongside neural data, we enable our model to focus on extracting dynamic visual information. This approach improves interpretability compared to models like MindVideo, which simultaneously predict both dynamic and static features. In contrast, our approach disentangles the specific contributions of dynamic components.
@@
In order to focus our evaluation on salient objects in a scene, we apply masking using FlowSAM [70] to track and identify salient objects. Consistent with claims in their paper, we found this method was effective in identifying salient objects.
@@
Figure 4: Side-by-side comparison of motion predicted with and without neural data. Motion predicted using brain data and a static reference image is qualitatively both plausible and aligned with the original video compared to motion predicted from only the reference image using Stable Video Diffusion [2]. Motion predicted from purely image features is plausible but not aligned with the original video. In this latter case, the model is hallucinating flows. (a) An example video where there are multiple plausible actions. (b) An example video where camera motion overrides the plausible motion. (c) An example video where there is ambiguity of stationary objects due to camera motion.
@@ 3.2 Visual Stimuli-To-Brain Encoding
      Self-supervised: VIP is a ResNet-50 based architecture trained on the large-scale human egocentric video dataset Ego4D. VC-1 is a vision transformer-based architecture trained on seven human egocentric video datasets, including Ego4D. For both models, we also test an encoder trained on features averaged over frames spanning the video (VIP AF and VC-1 AF [85, 12]).
@@
We train ridge regression models on visual stimuli features from different vision encoding models to predict fMRI voxel-wise responses.
@@ 4 Experiments
In this section, we describe the specific experiments we use to explore dynamic visual representations in the human brain. In 4.1, we describe the dataset and preprocessing used. In 4.2 we explore motion decoding (reconstruction) conditioned on initial image and brain information, and augment quantitative analysis with visualizations derived from a motion-conditioned diffusion model. Finally in 4.3 , we explore the performance of different visual backbones in predicting brain activations.
@@ 4.1 Dataset
In practice, we temporally downsample the flows to $c=3$ frames, selected in equal intervals over the corresponding 2 second fMRI brain scan, and aim to predict the $c$ quantized consecutive optical flow fields. We also spatially downsample the flow to a resolution of $32\times 32$ and quantize the flows into 40 clusters. We extract image features from the first frame using DINOv2. During test time, we take a weighted-sum over all the classes per pixel, following [77]. We train the model using cross entropy loss. We train all models on a single Nvidia Titan RTX with 24GB GPU RAM and 48GB CPU RAM for up to 12 hours. We illustrate our model’s pipeline in Figure 2.
fMRI captures the blood oxygen level dependent (BOLD) signal, which peaks approximately 4 seconds after stimulus presentation and slowly decays after, due to nature of the hemodynamic response. During preprocessing, the fMRI data is shifted to temporally align the peak of the response to the onset of its corresponding visual stimuli presentation. However, due to the long time scale of the fMRI response, fMRI scans 4 seconds or 2 TRs prior and after the current TR may have information shared with the current TR. In order to leverage all relevant neural responses, when predicting the motion in the video presented at TR=$i$, we concatenate fMRI data from TRs $[i-2,i-1,i,i+1,i+2]$. We generate the video frames conditioned on the predicted motion using DragNUWA [4].
@@ 4.2 Motion Decoding
Quantitative Flow Evaluation. First we report evaluations of the decoded optical flow vectors on the end point error on the test set. We quantitatively compare the models trained on the 3 participants against baselines of: MindVideo (Best) [3], our method leveraging the initial frame generated from MindVideo (Ours + MindVideo (Best)), and Stable Video Diffusion [2] that does not use neural data and only the initial frame in Figure 3. We extract the optical flows from the videos generated by MindVideo and Stable Video Diffusion using the same RAFT model we use to extract the ground truth flow. Since MindVideo and Stable Video Diffusion are generative models that can produce multiple plausible outputs, we generate 100 samples per video for MindVideo and 10 samples for Stable Video Diffusion (because the latter is able to condition on the true video’s first frame) and select the sample with the lowest end point error for fair comparison. We evaluate our method using the true initial frame from the video (Ours) as well as the first frame generated by MindVideo (Ours + MindVideo(Best)). Since the Stable Video Diffusion baseline (SVD (Best)) is not conditioned on neural data, it learns to predict plausible flows which do not necessarily align with the true flow. Indeed, we observe that our model trained on individual participant neural data performs significantly better on end point error as compared to SVD (Best) (end point error for individual participants – S1, S2, S3 – respectively, S1: 0.543, S2: 0.572, S3: 0.634 vs. 1.192, all $p\ll 0.001$ using paired $t$-tests with Holm-Bonferroni correction).
Reasoning With and Without fMRI. We present examples and explain why our models trained on neural data outperform the model without brain conditioning, Stable Video Diffusion (Figure  4).
First, there exists ambiguity of action given static images - a frame may contain multiple possible actions. For instance, in Figure 4a we show an eagle flying. Given the first frame, the eagle could be plausibly either flapping its wings or gliding. The model trained on neural data is able to identify that the eagle is gliding. However, the model without brain conditionining incorrectly predicts that the eagle is flapping its wings.
@@
Figure 5: Static image animation results. Two examples of optical flow predicted from fMRI brain activity (“Our Flow”) and the ground truth initial frame (gray box) compared against the ground truth (“GT”) flow. In addition, beneath the ground truth video frames we show our results for animating the initial frame by combining the brain conditioned motion prediction with DragNUWA [4] (“Our Video”).
Figure 6: Extending the method for video decoding. We demonstrate video decoding from fMRI by leveraging the initial frame generated by MindVideo [3] box). The ground truth flow (bottom-left) aligns more accurately with our predicted motion (bottom-right) than with the motion predicted by MindVideo (bottom-middle).
@@
Because we disentangle the modeling of static image and motion representations, we can extend our method to other works that decode image frames from fMRI. We demonstrate examples of extending our method to initial frames decoded from MindVideo in Figure 6 and find we decode motion more accurately (Figure 3). We provide additional video examples in the supplemental.

### v3

@@ Abstract
While computer vision models have made incredible strides in static image recognition, they still do not match human performance in tasks that require the understanding of complex, dynamic motion. This is notably true for real-world scenarios where embodied agents face complex and motion-rich environments. Our approach, BrainNRDS (Neural Representations of Dynamic Stimuli), leverages state-of-the-art video diffusion models to decouple static image representation from motion generation, enabling us to utilize fMRI brain activity for a deeper understanding of human responses to dynamic visual stimuli. Conversely, we also demonstrate that information about the brain’s representation of motion can enhance the prediction of optical flow in artificial systems. Our novel approach leads to four main findings: (1) Visual motion, represented as fine-grained, object-level resolution optical flow, can be decoded from brain activity generated by participants viewing video stimuli; (2) Video encoders outperform image-based models in predicting video-driven brain activity; (3) Brain-decoded motion signals enable realistic video reanimation based only on the initial frame of the video; and (4) We extend prior work to achieve full video decoding from video-driven brain activity. BrainNRDS advances our understanding of how the brain represents spatial and temporal information in dynamic visual scenes. Our findings demonstrate the potential of combining brain imaging with video diffusion models for developing more robust and biologically-inspired computer vision systems. We show additional decoding and encoding examples on this site: https://brain-nrds.github.io/.
@@ 1 Introduction
Figure 1: Encoding and decoding video motion using brain activity. (a) fMRI brain activity can be predicted (using a Pearson coefficient) using off-the-shelf video encoders (VideoMAE [1]) extracted from the viewed video. In the converse direction, we can generate video by decoding brain activity. (b) Many existing video diffusion models (e.g., SVD [2]) generate a video by animating an initial frame. (c) This suggests that“brain-to-video" generation can be achieved by fine-tuning diffusers to condition on fMRI input (e.g., MindVideo [3]). (d) In our approach we explicitly decouple the task of image and motion generation from brain activity. Given an initial video frame (which could be decoded from brain activity as in (c)) and fMRI input, we train a network to predict optical flow. We then animate the initial frame by feeding the predicted flow into an off-the-shelf motion-conditioned diffusion model (e.g., DragNUWA[4]). Our disentangled pipeline produces more accurate brain-conditioned motion decodings than either (b) or (c). (e) Ground truth optical flow and video.
@@
Human vision excels at understanding complex, dynamic scenes, while computer vision offers powerful tools for data processing and analysis. This work explores how insights from human brain activity, specifically in response to dynamic visual stimuli, can enhance the ability of computer vision systems to understand complex motion. Critically, as inherently embodied systems, our brains have evolved to process and represent the daily experience of naturally motion-rich environments in ways that are both semantically grounded and behaviorally-relevant. For example, a video of a person walking affords information not only about the visual features of the walker, but also about motion patterns, context, intent, and relationships between the different elements in the scene (e.g., sudden increase in gait implies urgency). This interplay between spatial and temporal information is impossible to capture from static images; accordingly, humans excel at inferring such high-level information from scene dynamics [5].
@@
Recent advances have improved our understanding of how visual information is represented in the brain through computer vision models [6, 7, 8], yet a significant gap persists in explaining how the brain processes dynamic visual stimuli. While some studies have started to explore the neural representation of dynamic visual stimuli [3, 9, 10], there remains substantial potential to further harness modern computer vision techniques, especially video-based models. There has also been little work demonstrating how insights from human vision can be leveraged to build better computer vision systems. Here we address both challenges, demonstrating that state-of-the-art video foundation models can be used to robustly predict human brain responses to dynamic visual inputs and that brain-decoded motion signals can help to reanimate videos based only on the first frame of the video. Critically, our framework outperforms prior work for both problems, achieving lower end point error (EPE) for motion decoding based on brain responses as compared to [3] and better motion prediction based on initial static frames with brain data than without brain data.
@@
In this paper, we present a framework for analyzing dynamic visual stimuli that explicitly separates the modeling of static and motion-related image representations as they are processed in the human brain: Neural Representations of Dynamic Stimuli (“BrainNRDS"). By decoupling these elements, BrainNRDS offers a targeted approach to understanding the distinct neural pathways used to process and represent dynamic visual information in real-world scenes. Better specified models of dynamic visual processing are also critical for developing effective models of ecologically critical domains such as social interaction [5], action representation, and dynamic causal structure [11] – none of which are easily extracted from static images.
@@
Contributions. First, we demonstrate that visual motion information in the form of optical flow can be predicted (or decoded) from brain activity as measured by fMRI. Second, we show that this predicted motion can be used to realistically animate static images using a motion-conditioned video diffusion model in which the motion is driven by brain activity. Moreover, we demonstrate that brain activity can produce more reliable animations than image-conditioned video diffusion models trained on massively-large datasets. Third, we apply BrainNRDS to existing methods that decode the viewed video by using their initial frame to do full video decoding from video-driven brain activity and find that we more accurately predict the viewed motion than existing methods. Finally, we show prediction in the reverse direction: recent video encoders (e.g. VideoMAE and VC-1 [1, 12]) more effectively predict fMRI brain activity from videos compared to image encoders. Overall these results demonstrate the effectiveness of BrainNRDS as a novel, extensible framework for interpreting how the human brain processes dynamic visual information.
@@
Larger Objectives. Our project also has several larger objectives. First, developing high-performing models for motion decoding based on neural data (Figure 1d) is an important precursor to building effective encoding models [13]. Second, the finding that VideoMAE is able to accurately predict brain responses (Figure 1a) is an important step towards building more general and informative video encoding models for brain prediction [7]). Encoding models are also important components in comparisons between model-based representations and empirical neural data, which can reveal key insights into which visual features the brain prioritizes and how they align with computational processing in interpretable artificial neural networks [14]. Third, we apply BrainNRDS to existing methods which decode the viewed video to do full video decoding from video-driven brain activity [3] and find that we more accurately predict the viewed motion than existing methods. Fourth, although comparison points are limited due to the underexploration of this research area, our finding that neural data improves video prediction performance better than other extant models is a simple demonstration of the point that latent information encoded in brain data may be useful for improving the performance of computer vision systems [15]. This approach may prove to be especially powerful for domains that are inherently anchored in human behavior, for example, dynamic social interactions.
@@ 2 Related Work
Static Visual Stimuli Decoding from fMRI. Reconstructing visual images from human brain activity, such as that measured by functional magnetic resonance imaging (fMRI), has significantly advanced through the use of deep learning models trained on natural image datasets [6, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 8, 28, 29, 30] and methods that leverage generative models to decode brain representations [31, 32, 3, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57]. This progress is largely due to the availability of large, publicly accessible fMRI datasets with well-annotated stimuli, such as BOLD5000 [58] and the Natural Scenes Dataset (NSD) [59], along with multimodal generative models that have been pre-trained on very large datasets, such as Stable Diffusion [60].
@@
Our fMRI-based approach offers a whole-brain perspective compared to single-neuron recordings, which are limited in spatial coverage [64, 65, 66, 61, 62, 67]. Despite the challenges of fMRI’s slow temporal resolution, our results demonstrate successful decoding of motion information from complex, naturalistic stimuli at the object level. Thus, fMRI, despite its limitations, can enable valuable insights into the neural representation of dynamic information.
Figure 2: BrainNRDS pipeline for motion decoding and video generation. (a) BrainNRDS takes in neural data and image features from DINOv2 [68], extracted from the initial frame to predict consecutive future dense optical flow fields. Salient objects are masked out using FlowSAM [69] to obtain the masked object flow. Snowflakes=frozen; flames=actively trained. (b) The initial frame is realistically animated using our predicted motion and a motion-conditioned video diffusion model, DragNUWA [4].
Concurrent to motion decoding research, there is a parallel branch of research focusing on decoding the raw video from fMRI activity [70, 71, 3, 72, 73, 74]. Methods like MindVideo and NeuroClips [3, 73] train to align temporal semantics from fMRI and generate frames across time. Notably, our approach explicitly disentangles the representations of motion information (represented as optical flow) from static images in fMRI signals. This allows us to differentiate our present work from previous work by enabling interpretability. Specifically, we isolate and focus on decoding motion (as optical flow) from fMRI signals, as opposed to more difficult to interpret mixture of static image features and motion. This disentanglement allows us to learn an explicit representation of motion which can be directly compared to groundtruth and be more directly related to spatially localized neural mechanisms, as opposed to an entangled mixture of static image and dynamic features. In contrast, NeuroClips [73] decodes blurred video as a proxy for motion. Comparing the two approaches, blurred video preserves rough scene composition, while optical flow explicitly encodes motion magnitude and direction, offering links to neural motion processing. At the same time, we ensure temporal consistency by integrating predicted optical flow with a motion-conditioned video diffusion model.
@@ 3.1 Motion Prediction
We extract the information relevant to visual motion in order to predict the optical flow in videos observed by participants. To enhance interpretability, we’ve architecturally separated static image features from dynamic features in BrainNRDS. By conditioning the model on the initial frame alongside neural data, we enable our model to focus on extracting dynamic visual information. This approach improves interpretability compared to models like MindVideo, which simultaneously predict both dynamic and static features. In contrast, our approach disentangles the specific contributions of dynamic components.
@@
In order to focus our evaluation on salient objects in a scene, we apply masking using FlowSAM [69] to track and identify salient objects.
@@
Figure 4: Motion predicted with and without neural data. Motion predicted using brain data and a static reference image is qualitatively both plausible and aligned with the original video compared to motion predicted from only the reference image using Stable Video Diffusion [2]. Motion predicted from purely image features is plausible but not aligned with the original video. In this latter case, the model is hallucinating flows. (a) An example video where there are multiple plausible actions. (b) An example video where camera motion overrides the plausible motion. (c) An example video where there is ambiguity of stationary objects due to camera motion.
Method Video-based Frame-based
[…]
MindVideo 0.742$\pm$0.006 0.879$\pm$0.004 0.171$\pm$0.02
Ours end-to-end 0.769$\pm$0.006 0.896$\pm$0.003 0.214$\pm$0.01
Subject 1 0.771$\pm$0.006 0.896$\pm$0.003 0.218$\pm$0.01
Subject 2 0.774$\pm$0.006 0.896$\pm$0.003 0.2$\pm$0.01
Subject 3 0.762$\pm$0.006 0.894$\pm$0.003 0.223$\pm$0.01
Table 1: Video generation quantitative results. Our end-to-end method uses the initial frame generated from fMRI by MindVideo and generates the video solely from fMRI without access to GT frames. We report semantic fidelity metrics and SSIM on both video and frame level. Results are significant ($p<$0.05 using two-sample t-test) compared to MindVideo.
@@ 3.2 Visual Stimuli-To-Brain Encoding
      Self-supervised: VIP is a ResNet-50 based architecture trained on Ego4D. VC-1 is a vision transformer-based architecture trained on seven human egocentric video datasets, including Ego4D. For both models, we also test an encoder trained on features averaged over frames spanning the video (VIP AF and VC-1 AF [84, 12]).
@@
We train ridge regression models on visual features from different encoding models to predict voxel-wise responses.
@@ 4 Experiments
In 4.1, we describe the dataset and preprocessing used. In 4.2 we explore motion decoding (reconstruction) conditioned on initial image and brain information using BrainNRDS, and augment quantitative analysis with visualizations derived from a motion-conditioned diffusion model. In 4.3, we explore the performance of different visual backbones in predicting brain activations.
@@ 4.1 Dataset
In practice, we temporally downsample the flows to $c=3$ frames, selected in equal intervals over the corresponding 2 second fMRI brain scan, and aim to predict the $c$ quantized consecutive optical flow fields. We also spatially downsample the flow to a resolution of $32\times 32$ and quantize the flows into 40 clusters. We extract image features from the first frame using DINOv2. During test time, we take a weighted-sum over all the classes per pixel, following [76]. We train the model using cross entropy loss. We illustrate our model’s pipeline in Figure 2.
fMRI signal peaks approximately 4 seconds after stimulus presentation and slowly decays due to the hemodynamic response. In preprocessing, the fMRI data is shifted to temporally align the peak of the response to the onset of its corresponding stimulus presentation. However, fMRI scans 4 seconds(2 TRs) prior and after the current TR may share information with the current TR. Consequently, when predicting the motion in the video presented at TR=$i$, we concatenate data from TRs $[i-2,i-1,i,i+1,i+2]$. Video frames are generated conditioned on the predicted motion using DragNUWA [4].
@@ 4.2 Motion Decoding
Architecture. BrainNRDS predicts motion labels by processing fMRI and image features. fMRI data is processed through an MLP then spatially broadcasted and concatenated over the image features. The concatenated features pass through three residual 1x1 convolutional blocks with dropout, while global features are extracted using AdaptiveAvgPool2d and concatenated back into the spatial features. A final 1x1 convolution classifies each spatial patch into the motion codebook.
Quantitative Flow Evaluation. We report evaluations of the decoded optical flow vectors on the end point error on the test set. We quantitatively compare the models trained on the 3 participants against baselines of: MindVideo (Best) [3], BrainNRDS leveraging the initial frame generated from MindVideo (Ours + MindVideo (Best)), and Stable Video Diffusion (SVD) [2] that does not use neural data and only the initial frame in Figure 3. We extract the optical flows from the videos generated by MindVideo and SVD using the same RAFT model we use to extract the ground truth flow. Since MindVideo and SVD are generative models that can produce multiple plausible outputs, we generate 100 samples per video for MindVideo and 10 samples for SVD (because the latter is able to condition on the true video’s first frame) and select the sample with the lowest end point error for fair comparison. We evaluate BrainNRDS using the true initial frame from the video (Ours) as well as the first frame generated by MindVideo (Ours + MindVideo(Best)). Since the SVD baseline (SVD (Best)) is not conditioned on neural data, it learns to predict plausible flows which do not necessarily align with the true flow. Indeed, we observe that our model trained on individual participant neural data performs significantly better on end point error compared to SVD (Best) (end point error for individual participants – S1, S2, S3 – respectively, S1: 0.543, S2: 0.572, S3: 0.634 vs. 1.192, all $p\ll 0.001$ using paired $t$-tests with Holm-Bonferroni correction).
Reasoning With and Without fMRI. We present examples and explain why models trained on neural data outperform SVD (Figure  4). First, there exists ambiguity of action given static images - a frame may contain multiple possible actions. For instance, in Figure 4a we show an eagle flying. Given the first frame, the eagle could be plausibly either flapping its wings or gliding. The model trained on neural data is able to identify that the eagle is gliding. However, the model without brain conditionining incorrectly predicts that the eagle is flapping its wings.
@@
Figure 5: Static image animation results. Optical flow predicted from fMRI brain activity (“Our Flow”) and the ground truth initial frame (gray box) compared against the ground truth (“GT”) flow. Beneath the ground truth video frames, we show our results for animating the initial frame by combining the brain conditioned motion prediction with DragNUWA [4] (“Our Video”).
Figure 6: Extending the method for video decoding. Video decoding from fMRI is accomplished by leveraging the initial frame generated by MindVideo [3] box). The ground truth flow (bottom-left) better aligns with our predicted motion (bottom-right) than with the motion predicted by MindVideo (bottom-middle).
@@
Because we disentangle the modeling of static image and motion representations, we can extend BrainNRDS to other works that decode image frames from fMRI. We demonstrate examples of extending our method to initial frames decoded from MindVideo in Figure 6 and find we decode motion more accurately (Figure 3). We provide additional video examples in the supplemental.
Generated videos are evaluated using initial frames decoded from MindVideo (Table 1). Video-level evaluation is computed using the cosine similarity between averaged VideoMAE features of the ground truth and generated videos. Frame-level evaluation is computed using the cosine similarity between CLIP features of corresponding frames. SSIM is also reported. BrainNRDS outperforms MindVideo across all metrics, demonstrating the advantage of separately modeling spatial and temporal information.